some cooking pans have a layer of cooper on the bottom.how many kilojoules are needed to raised rhe temperature of 125 degree Celsius to 325 degree celsius
1
Expert's answer
2018-02-07T13:27:08-0500
We have to use the equation to find heat energy:
Q=cm∆T,
where
Q is heat energy;
c is heat capacity (for copper c = 0.39 kJkg-1K-1);
m is a mass of the material, kg
∆T= T_2-T_1=325°C-125°C=220°C=220K
The difference between temperatures in degree Kelvin is the same that in degree Celsius.
Q=kJ/(kg×K)×kg×K=kJ
Q=0.39×m×220=85.8m
To find how many kJ are needed we have to know the mass of the copper layer in kg.
Answer:
Q = 85.8m (kJ)